OpenStudy (anonymous):

Factor the following completely. Look first for the greatest common factor: 2x^4 − 24x^3 + 54x^2

OpenStudy (anonymous):

2x^2(x^2 - 12x + 27)

OpenStudy (anonymous):

2x^2(x-9)(x-3)

OpenStudy (anonymous):

Is that the answer?

OpenStudy (anonymous):

those are the factors of the equation. do you want the "x=" things?

OpenStudy (anonymous):

I think so, it says factor completely.

OpenStudy (anonymous):

okay well set each of those 3 factors equal to 0 2x^2=0 divide both sides by 2 u get x^2=0 then square root both sides and u get x=0 then x-9=0 add 9 to both sides and u get x=9 and lastly x-3=0 add 3 to both sides and u get x=3 ur answers are x=0,x=3, and x=9